OpenStudy (karolynxx):

Muriel says she has written a system of two linear equations that has an infinite number of solutions. One of the equations of the system is 3y = 2x – 9. Which could be the other equation? 2y = x – 4.5 y = x – 3 6y = 6x – 27 y = x – 4.5

OpenStudy (anonymous):

((1/3y)) ((1/6y))/((1/2y 3/4y)) is 4/27.

OpenStudy (karolynxx):

so C?

OpenStudy (anonymous):

2y = x – 4.5 y = x – 3 6y = 6x – 27 y = x – 4.5

OpenStudy (anonymous):

no, i was wrong d

OpenStudy (karolynxx):

are you sure?

OpenStudy (anonymous):

thats what i saw on weegy.com, so dont blame me if im wrong. tag a couple people u know